OpenStudy (anonymous):

solve using elimination method 5x+5y=-7 7x-2y=13

OpenStudy (anonymous):

ok, lets eliminate y, and find x first: to eliminate y lets multiply 1st eq by 2 and 2nd eq by -5, to make both Ys cancel each other out once subtracted, so we will have: 10x+10y=-70 -35x+10y=-65 so now we have to subtract coresponding members in both eqs: 45x+0=-5 45x=-5 now divide both sides by a 45 to find x x=-5/45=-0.111

OpenStudy (anonymous):

to find y just substitute it into any of the equations: 5x+5y=-7 5*-0.11111+5y=-7 5y=-7+0.5555 5y=-6.44445 y=-6.44445/5 y = -1.28889 y = -1.289

OpenStudy (anonymous):

now lets see if I was correct, by substituting the values into any eq: 5x+5y=-7 5*-0.1111+5*-1.289= =-7 so I was right and the answers are correct

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

that doesn't work for the other equation...you messed up somewhere on top

OpenStudy (anonymous):

:/ sorry if I did

OpenStudy (anonymous):

but works for me

OpenStudy (anonymous):

you didn't cancel out y because you had positive 10y on both equations instead of one positive and one negative so that only gave you the right answer for one equation

OpenStudy (anonymous):

ok, so how do I do it?

OpenStudy (anonymous):

do the first equation the same but instead of multiplying second equation by -5 multiply it by 5 so you get 10x + 10y = -14 35x - 10y = 65 this way y really is eliminated then you have 45x = 51 so x = 51/45 which is approximately 1.1333.... plug into either equation and solve 7(1.1333) - 2y = 13 or 7.9331 -2y = 13 after solving for y you get y = -2.53345 now check those two values in both equations

OpenStudy (anonymous):

ok, thanx mate